### v1

# Jet production at electron-positron colliders at next-to-next-to-next-to-leading order in QCD

Abstract
We present the first application of antenna subtraction at next-to-next-to-next-to-leading order (N3LO) in QCD by computing fully differential predictions for two-jet production at electron-positron colliders. We illustrate the structure of the infrared counterterms and provide results for the N3LO correction to the two-jet production rate and to the leading-jet energy. Our work constitutes the first direct calculation of jet production at electron-positron colliders at N3LO and represents the first step in tackling arbitrary processes with jets at this perturbative order.
preprint: IPPP/25/27, ZU-TH 36/25
Introduction—Experiments at present and future colliders plan to deliver an unprecedented amount of data and significantly increase our sensitivity to new physics if equally precise Standard Model predictions are available. In high-energy collisions, the effects of QCD are dominant and their evaluation at higher perturbative orders in the strong coupling $\alpha_{s}$ requires techniques for the identification and removal of infrared (IR) divergences from the constituent scattering amplitudes, known as infrared subtraction schemes.
(a) triple-real
(b) double-real-virtual
(c) real-double-virtual
(d) triple-virtual
Figure 1: Representative Feynman diagrams (before interference with an appropriate tree-level diagram) for each type of radiative correction to the process $e^{+}e^{-}\to 2\,$jets at $\mathcal{O}(\alpha_{s}^{3})$. From left to right: $\text{d}\sigma^{(3)}_{5}$, $\text{d}\sigma^{(3)}_{4}$, $\text{d}\sigma^{(3)}_{3}$, and $\text{d}\sigma^{(3)}_{2}$.
General approaches have been developed for next-to-leading order (NLO) calculations [1, 2], followed by several strategies at next-to-next-to-leading order (NNLO) [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and ongoing efforts to include higher-multiplicity final states and to generalize the current formalisms [17, 18, 19, 20, 21]. NNLO-accurate predictions are nowadays available for several Standard Model processes, see [22] for a review. N3LO accuracy has recently been reached for color-singlet production and decay processes for both inclusive [23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and differential [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] observables, relying either on $q_{T}$-slicing [7] or the Projection-to-Born [11] method. Neither of these approaches is applicable to arbitrary processes and observables, especially in the presence of final-state jets, making the development of a general N3LO subtraction scheme highly desirable.
The antenna subtraction method, formulated at NLO in [43, 44] and extended to NNLO in [3, 5] is one of the most flexible local subtraction schemes with the ability to handle jets at both lepton and hadron colliders. Its fundamental ingredients are antenna functions [45, 46, 47], which encode the singular behavior of real-emission matrix elements and can be analytically integrated over the phase space of the unresolved radiation in order to cancel the explicit IR singularities of virtual corrections.
In this Letter, we present the first application of antenna subtraction to fully differential predictions for physical cross sections at N3LO. We focus on the process $e^{+}e^{-}\to jj+X$, which was computed at NNLO with the sector decomposition [48] and antenna subtraction [49] methods. In [50], the N3LO correction to the two-jet production rate was evaluated differentially in the Durham jet reconstruction algorithm parameter $y_{\text{cut}}$ [51, 52, 53, 54, 55] indirectly, i.e. by combining an NNLO calculation for three-jet production with the analytic result for the N3LO total inclusive cross section for $e^{+}e^{-}\to$ hadrons [56]. In this Letter, we evaluate the two-jet rate with a direct method that can in principle account for any experimental cuts. In addition, we determine the distribution of the leading-jet energy in two-jet events, previously unknown at $\mathcal{O}(\alpha_{s}^{3})$.
This work represents the first N3LO differential calculation obtained with a subtraction scheme that is not restricted to specific classes of processes and does not rely on the prior knowledge of inclusive quantities. Our calculation is therefore a crucial step in tackling more complicated processes with jets at high perturbative order [57] relevant to the physics goals of the upcoming High-Luminosity phase of the LHC [58]. Moreover, results at this level of precision are relevant for future QCD precision studies [59] at the next generation of electron-positron colliders, such as the FCC-ee [60, 61, 62] and CEPC [63, 64, 65].
Method—In the following, we outline the extension of antenna subtraction for final-state radiation to N3LO. The $\mathcal{O}(\alpha_{s}^{3})$ correction to two-jet production at electron-positron colliders receives contributions from final states with up to five partons. We denote the contribution to the N3LO coefficient from a configuration with $n$ final-state partons by $\text{d}\sigma^{(3)}_{n}$. The triple-real contribution $\text{d}\sigma^{(3)}_{5}$ is given by three real emissions on top of the Born-level process, the double-real-virtual contribution $\text{d}\sigma^{(3)}_{4}$ by two real emissions at one loop, the real-double-virtual contribution $\text{d}\sigma^{(3)}_{3}$ by one real emission at two loops and finally the triple-virtual contribution $\text{d}\sigma^{(3)}_{2}$ by the three-loop correction. Examples of Feynman diagrams contributing to different types of N3LO corrections are shown in Figure 1.
The individual contributions contain explicit IR singularities due to loop corrections ($\text{d}\sigma^{(3)}_{2}$), implicit singularities arising when the external momenta approach soft and collinear limits ($\text{d}\sigma^{(3)}_{5}$), or a mixture of both ($\text{d}\sigma^{(3)}_{3}$ and $\text{d}\sigma^{(3)}_{4}$). These divergences cancel when virtual and real corrections are combined in physical predictions [66, 67], but each individual term is numerically integrable over the associated phase space only after introducing dedicated subtraction terms. We denote by $\text{d}\sigma^{(3)}_{n,\text{finite}}$ the individual contributions after subtraction, free from any type of IR singularity. The definition of $\text{d}\sigma^{(k)}_{n,\text{finite}}$ up to NNLO in the context of antenna subtraction was described in [3, 5]. At N3LO, we write schematically
$\text{d}\sigma_{5,\text{finite}}^{(3)}$ $=\text{d}\sigma_{5}^{(3)}$ $-\text{d}\sigma^{(3)}_{5,S_{1}}$ $-\text{d}\sigma^{(3)}_{5,S_{2}}$ $-\text{d}\sigma^{(3)}_{5,S_{3}}\,,$ (1)
$\text{d}\sigma_{4,\text{finite}}^{(3)}$ $=\text{d}\sigma_{4}^{(3)}$ $-\text{d}\sigma^{(3)}_{4,V_{1}}$ $-\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}$ $-\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}\,,$ (2)
$\text{d}\sigma_{3,\text{finite}}^{(3)}$ $=\text{d}\sigma_{3}^{(3)}$ $-\text{d}\sigma^{(3)}_{3,V_{2}}$ $-\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}\,,$ (3)
$\text{d}\sigma_{2,\text{finite}}^{(3)}$ $=\text{d}\sigma_{2}^{(3)}$ $-\text{d}\sigma^{(3)}_{2,V_{3}}\,,$ (4)
where $\text{d}\sigma^{(3)}_{n,V_{\ell}S_{m}}$ represents a counterterm targeting the divergent behavior of $m$ unresolved partons as well as the explicit IR poles in a $\ell$-loop matrix element.
The following relations between the subtraction terms ensure the cancellation of IR singularities across different phase-space multiplicities:
$\text{d}\sigma^{(3)}_{4,V_{1}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{5,S_{1}}\,,$ (5)
$\text{d}\sigma^{(3)}_{3,V_{2}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}-\int_{2}\text{d}\sigma^{(3)}_{5,S_{2}}\,,$ (6)
$\text{d}\sigma^{(3)}_{2,V_{3}}$ $=$ $-\int_{1}\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}-\int_{2}\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}-\int_{3}\text{d}\sigma^{(3)}_{5,S_{3}}\,,$ (7)
where $\int_{m}$ indicates inclusive integration over the phase space of $m$ unresolved emissions. As a consequence of the relations (5)–(7), the net contribution of the subtraction terms vanishes in the sum of (1)–(4), restoring the correct physical result
$\sum_{n=2}^{5}\int_{n}\text{d}\sigma^{(3)}_{n,\text{finite}}=\sum_{n=2}^{5}\int_{n}\text{d}\sigma^{(3)}_{n}\,,$ (8)
with the crucial difference that each term in the sum on the left-hand side is individually well-defined and computable with numerical techniques.
The subtraction terms $\text{d}\sigma^{(3)}_{5,S_{1}}$, $\text{d}\sigma^{(3)}_{5,S_{2}}$, $\text{d}\sigma^{(3)}_{4,V_{1}}$, $\text{d}\sigma^{(3)}_{4,V_{1}S_{1}}$ and $\text{d}\sigma^{(3)}_{3,V_{2}}$ are analogous to those required for an NNLO calculation of three-jet production in electron-positron annihilation [68, 69]. The subtraction terms $\text{d}\sigma^{(3)}_{5,S_{3}}$, $\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}$, $\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}$ and $\text{d}\sigma^{(3)}_{2,V_{3}}$ specifically target N3LO infrared singularities and have been constructed for the first time for the present calculation.
The subtraction terms are built from antenna functions with the following notation: $X_{n}^{\ell}$ is a $n$-parton $\ell$-loop antenna function intended to subtract the singular behavior of $n-2$ unresolved partons between a pair of hard radiators at $\ell$ loops. $X$ is a placeholder for a particular partonic configuration [3]. The quark-antiquark antenna functions relevant for our process are obtained from color-ordered matrix elements for the decay of a photon to $n$ partons [45, 70], denoted by $M_{n,\gamma^{*}}^{\ell}$. The integrated counterpart of an antenna function, $\mathcal{X}_{n}^{\ell}$, is obtained by analytic integration over the inclusive phase space of the $n-2$ unresolved emissions. All the N3LO final-state antenna functions were recently integrated in [70, 71, 72].
The infrared limits appearing at N3LO cannot in general be addressed by a combination of NLO and NNLO ingredients, and three new types antenna functions must be defined: five-parton tree-level
$X_{5}^{0}$ $=N_{5}\,\dfrac{M_{5,\gamma^{*}}^{0}}{M_{2,\gamma^{*}}^{0}}\,,$ (9)
four-parton one-loop
$X_{4}^{1}$ $=N_{4}\,\dfrac{M_{4,\gamma^{*}}^{1}}{M_{2,\gamma^{*}}^{0}}-X_{4}^{0}\dfrac{M_{2,\gamma^{*}}^{1}}{M_{2,\gamma^{*}}^{0}}\,,$ (10)
and three-parton two-loop
$X_{3}^{2}$ $=N_{3}\dfrac{M_{3,\gamma^{*}}^{2}}{M_{2,\gamma^{*}}^{0}}-X_{3}^{1}\dfrac{M_{2,\gamma^{*}}^{1}}{M_{2,\gamma^{*}}^{0}}-X_{3}^{0}\dfrac{M_{2,\gamma^{*}}^{2}}{M_{2,\gamma^{*}}^{0}}\,,$ (11)
where $N_{i}$ indicates a numerical symmetry factor and the two-body decay matrix element $M_{2,\gamma^{*}}^{0}$ is a normalization factor. In the case of the loop antenna functions $X_{4}^{1}$ and $X_{3}^{2}$, we remove the singular behavior that can be captured by antenna functions of lower loop order, naturally extending the definition of three-parton one-loop antenna functions at NNLO [3].
The $X_{5}^{0}$ antenna functions are used in the subtraction term $\text{d}\sigma^{(3)}_{5,S_{3}}$ to remove triple-real unresolved configurations. The singular behavior of the double-real-virtual correction is addressed with the functions $X_{4}^{1}$ within the term $\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}$, while $X_{3}^{2}$ are used in $\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}$ to subtract the soft and collinear divergences of the real-double-virtual correction. Finally, the term $\text{d}\sigma^{(3)}_{2,V_{3}}$ collects the integrated counterparts of all the N3LO antenna functions ($\mathcal{X}_{5}^{0}$, $\mathcal{X}_{4}^{1}$, $\mathcal{X}_{3}^{2}$) and removes the explicit IR singularities of the triple-virtual correction. Throughout, we work with spin-averaged antenna functions and we fully remove azimuthal correlations in collinear limits using the angular averaging routines described in [73], extended to handle triple-unresolved configurations.
After the removal of IR poles, the construction of $\text{d}\sigma^{(3)}_{3,V_{2}S_{1}}$ and $\text{d}\sigma^{(3)}_{4,V_{1}S_{2}}$ closely follows the logic of single-real and double-real infrared subtraction, already considered up to NNLO. A subtlety arises when a one-loop reduced matrix element and a one-loop unresolved factor appear in a product. To eliminate the need for higher orders in the dimensional regulator $\epsilon$ in one-loop matrix elements, we consistently work with IR finite remainders.
Let us elaborate on a crucial aspect of a N3LO local subtraction scheme—the removal of triple-unresolved limits. We start by recalling the structure of single-unresolved and double-unresolved subtraction terms:
$\text{d}\sigma_{5,S_{1}}^{(3)}$ $=$ $X_{3}^{0}\otimes M_{4}^{0}J_{2}^{(4)},$ (12)
[…]
$+\sum_{acc}(X_{3}^{0}\otimes X_{3}^{0}+\mathcal{S}\otimes X_{3}^{0})\Big{]}$ (13b)
$\otimes M_{3}^{0}J_{2}^{(3)},$
where $M_{n}^{0}$ represents a generic tree-level color-ordered matrix element for $e^{+}e^{-}\to n$ partons and the shorthand symbol $\otimes$ encodes a momentum mapping from a higher-multiplicity to a lower-multiplicity phase space, in order to absorb unresolved emissions. For each term, a sum over external partons and color correlations is understood. The jet function $J_{n_{j}}^{(n_{p})}$ reconstructs $n_{j}$ jets from $n_{p}$ final-state partons. The sums in (13a) and (13b) run over configurations where the two unresolved emissions are respectively color-connected ($cc$) and almost color-connected ($acc$[3, 68].
The singular behavior of color-connected emissions can only be captured by dedicated $X_{4}^{0}$ antenna functions. To prevent double-counting of IR limits already addressed by (12), we introduce the term $X_{3}^{0}\otimes X_{3}^{0}$ in (13a), which subtracts the single-unresolved singularities from the $X_{4}^{0}$. In the second term of (13b), $\mathcal{S}$ represents a combination of eikonal factors required to remove soft singularities at large angle [74, 68]. Finally, the $X_{3}^{0}\otimes X_{3}^{0}$ and $\mathcal{S}\otimes X_{3}^{0}$ structures present in $\text{d}\sigma_{5,S_{2}}^{(3)}$ cancel against $\text{d}\sigma_{5,S_{1}}^{(3)}$ in (12) when the reduced matrix element $M_{4}^{0}$ develops an additional singularity on top of that already captured by the $X_{3}^{0}$ antenna function. Together, $\text{d}\sigma_{5,S_{1}}^{(3)}+\text{d}\sigma_{5,S_{2}}^{(3)}$ subtract all the single- and double-unresolved behavior of the five-parton tree-level matrix element fully differentially across the phase space.
The triple-real subtraction terms have an analogous structure:
$\text{d}\sigma_{5,S_{3}}^{(3)}=\Big{[}$ $\phantom{-}X_{5}^{0}$ (14a)
$-X_{3}^{0}\otimes X_{4}^{0}$ (14b)
$-\sum_{cc}(X_{4}^{0}-X_{3}^{0}\otimes X_{3}^{0})\otimes X_{3}^{0}$ (14c)
$-\sum_{acc}(X_{3}^{0}\otimes X_{3}^{0}+\mathcal{S}\otimes X_{3}^{0})\otimes X_{3}^{0}\Big{]}$ (14d)
$\phantom{-}\otimes M_{2}^{0}J_{2}^{(2)}$
Here the $X_{5}^{0}$ antenna functions address color-connected triple-unresolved emissions. In particular, they contain the singular behavior in triple-soft [75, 76, 77] and quadruple-collinear [78, 79] limits. A $5\to 2$ antenna momentum mapping [80] is used to relate the triple-real and the fully resolved phase space. The remaining terms are needed to remove the single- and double-unresolved limits of the $X_{5}^{0}$, which are already subtracted by $\text{d}\sigma^{(3)}_{5,S_{1}}+\text{d}\sigma^{(3)}_{5,S_{2}}$. The five-particle antenna functions mimic the full matrix element in unresolved limits and factorize onto lower-multiplicity antenna functions as follows:
[…]
Once again, the over-subtraction of single-unresolved singularities is prevented by removing the iterated $X_{3}^{0}\otimes X_{3}^{0}$ contribution from the $X_{4}^{0}$ using the second term in (14c).
The iterated structures in $\text{d}\sigma_{5,S_{3}}^{(3)}$ also remove the singular behavior of $\text{d}\sigma_{5,S_{1}}^{(3)}$ and $\text{d}\sigma_{5,S_{2}}^{(3)}$ in triple-unresolved limits. In particular, (14b) cancels the additional double-unresolved singularities of $M_{4}^{0}$ in (12), while (14b) and (14c) subtract the single-unresolved behavior of $M_{3}^{0}$ originating from the terms (13a) and (13b), respectively. Finally, note that in $\text{d}\sigma_{5,S_{3}}^{(3)}$, there is no analogue of the term (13b) in $\text{d}\sigma_{5,S_{2}}^{(3)}$. This is due to the fact that with up to five partons in the final state, the three unresolved emissions must occur between the same pair of hard radiators, and they are therefore always color-connected. For higher-multiplicity processes, such as three-jet production at N3LO, additional terms will be required to subtract singularities due to three unresolved partons between three or more hard radiators; for a discussion at NNLO see [18].
Calculation—Our calculation was performed with a version of the NNLOJET Monte Carlo framework [81] extended for an N3LO calculation. For the N3LO antenna functions $X_{n}^{\ell}$, we adapted the implementation of the relevant matrix elements in EERAD3 [82]. The necessary helicity amplitudes for $e^{+}e^{-}\to\text{partons}$ are available in NNLOJET as part of the calculation of three-jet production at NNLO [83]. The five-parton tree-level amplitudes used in our calculation were computed in [84, 85, 86], the four-parton one-loop amplitudes in [87, 88, 89] and the three-parton two-loop amplitudes in [90, 91]. Finally, the three-loop quark form factor was computed in [92, 93, 94].
The numerical evaluation of these matrix elements in double-unresolved configurations at one loop and single-unresolved limits at two loops required careful treatment to achieve the necessary accuracy. In double-unresolved limits of one-loop matrix elements, we implemented a dynamical switch from double to quadruple precision for floating point operations during the evaluation of the $X_{4}^{1}$ antenna functions at phase-space points in the unresolved regions. The switch is triggered when a two-particle invariant is lower than a given threshold, $s_{ij}/s<10^{-5}$, where $s$ is the center-of-mass energy squared. The two-loop quantities are expressed in terms of two-dimensional harmonic polylogarithms [95]. These functions can be numerically evaluated using the library tdhpl [96], which uses a representation optimized to reduce the number of arithmetic operations for points in the bulk. For points with $s_{qg}/s<10^{-6}$ or $s_{\bar{q}g}/s<10^{-6}$, we substitute the amplitudes with their series expansions retaining divergent, constant and linear terms in the small invariant, obtained analytically using an algorithm [97] implemented in PolyLogTools [98].
To validate our construction of the antenna subtraction terms at N3LO, we verified that they cancel against the relevant matrix elements in all single-, double- and triple-unresolved limits at a satisfactory level for numerical integration with Monte Carlo techniques. The complete analytic cancellation of the $\epsilon$ poles of the virtual contributions is another check on the subtraction terms.
Figure 2: The two-jet production rate up to N3LO, differential in the Durham jet resolution parameter $y_{\text{cut}}$. Error bars indicate the Monte Carlo error.
Results—All our results are normalized by the respective Born-level cross section $\sigma_{0}$ for $e^{+}e^{-}\to q\bar{q}$. As a consequence, the electroweak coupling drops out and our results are valid for both $\gamma^{*}$ and $Z$ as the intermediate boson. We neglect the singlet contribution which is numerically small, as observed for example in [99, 100]. We set $\sqrt{s}=m_{Z}=91.2$ GeV and $\alpha_{s}(m_{Z})=0.118$. However, for the observables we consider, the dependence on the value of the energy scale only enters through the running of the strong coupling. The renormalization scale is chosen to be equal to the center-of-mass energy, $\mu_{r}=\sqrt{s}$, and theory uncertainties are estimated by varying it within $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$.
The first result of this Letter is the $\mathcal{O}(\alpha_{s}^{3})$ correction to the inclusive cross section for hadron production. For five light quark flavors, we find
$\sigma^{(3)}$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-105\pm 11)\,,$ (18)
where the error is due to Monte Carlo integration. This value is in agreement with the analytic result [56] given by $\approx-102.14$. We further compute the $\mathcal{O}(\alpha_{s}^{3})$ correction to the two-jet exclusive cross section at two different values of the Durham [51, 52, 53, 54, 55] jet reconstruction parameter:
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.1)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-455\pm 32)\,,$ (19)
$\sigma^{(3)}_{2j}(y_{\text{cut}}=0.01)$ $=\sigma_{0}\left(\frac{\alpha_{s}}{2\pi}\right)^{3}(-2554\pm 57)\,.$ (20)
We verified that both values are compatible with those obtained as the difference between the inclusive cross section and the three-jet production cross section at NNLO. The relative error is smaller for (20) because it is less affected by large cancellations occurring in more inclusive setups.
Figure 3: Distribution of the energy of the more energetic jet $E_{j,1}$ as a fraction of the center-of-mass energy $\sqrt{s}$ for events with two jets reconstructed with the Durham algorithm and $y_{\text{cut}}=0.1$ (left) or $y_{\text{cut}}=0.01$ (right). Error bars indicate the Monte Carlo error.
As the first example of a differential distribution, we compute the two-jet production rate up to $\mathcal{O}(\alpha_{s}^{3})$ differentially in the resolution parameter $y_{\text{cut}}$ of the Durham algorithm, shown in Figure 2. We also verified that we can reproduce the result of the indirect calculation presented in Figure 2 of [50] after accounting for a different value of $\alpha_{s}$.
As a second example, we consider the bin-integrated distribution of the energy of the more energetic jet in exclusive two-jet events (in the four-momentum recombination scheme), which was calculated at NNLO in [48] using the JADE [101] algorithm. We recovered the NNLO results and we computed for the first time the $\mathcal{O}(\alpha_{s}^{3})$ correction shown in Figure 3 for the Durham algorithm with $y_{\text{cut}}=0.1$ and $y_{\text{cut}}=0.01$. For $y_{\text{cut}}=0.1$, we observe very good perturbative convergence with percent-level N3LO corrections lying within the theoretical uncertainty of the previous order across the entire considered range. On the other hand, for $y_{\text{cut}}=0.01$, the NLO and NNLO corrections quickly vanish as the energy increases and, accordingly, the N3LO correction becomes comparatively large because the nominal perturbative accuracy in the tail of the distribution is lower. Indeed, a highly energetic jet must recoil against multiple partonic emissions, which are more likely to be clustered as three or more jets for a smaller value of $y_{\text{cut}}$. For both choices of $y_{\text{cut}}$, real radiation is suppressed close to $E_{j,1}=\sqrt{s}/2$, affecting the cancellation of IR singularities and calling for all-order resummation in order to obtain physical results, as discussed e.g. in [41].
Conclusions—We presented the first direct fully differential calculation of jet production in electron-positron annihilation at third order in QCD. We described the structure of the IR counterterms for the final-state radiation in $e^{+}e^{-}\to jj$ within antenna subtraction, including novel triple-unresolved limits. As an application, we computed the production rate of two jets and the distribution of the energy of the leading jet at N3LO at different values of the jet reconstruction parameter.
The future generalization of the presented N3LO antenna subtraction terms to higher-multiplicity final states and hadronic initial states will likely build on the definition of optimal antenna functions [102, 103, 104, 18] and recent steps towards the integration of N3LO antenna functions for initial-state radiation [105]. Our work is an important milestone in establishing a general method for computing N3LO predictions to processes involving final-state jets.
Acknowledgements.
Acknowledgments—We are grateful to Thomas Gehrmann, Nigel Glover and Aude Gehrmann-de Ridder for support and encouragement to pursue this work. We thank Elliot Fox for comments on the manuscript. XC is supported by the National Science Foundation of China (NSFC) with grant No.12475085 and No.12321005. PJ is supported by the European Research Council (ERC) under the European Union’s research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP). MM is supported by a Royal Society Newton International Fellowship (NIF/R1/232539). We acknowledge financial support, supercomputing resources and support from ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – and hosting entity, funded by European Union – NextGenerationEU. We are grateful to each other’s institutions and the Pauli Center for Theoretical Studies for hospitality.